OpenStudy (anonymous):

"Solve the equation. If necessary, round to the nearest hundredth. If there is no solution, write no solution. k^2-24k+4= -2 " Can someone help me solve this?

OpenStudy (misty1212):

HI!

OpenStudy (anonymous):

BYE!

OpenStudy (misty1212):

you want to add 2 to both sides

OpenStudy (misty1212):

start at \[k^2-24k+6=0\] then maybe factor it

OpenStudy (misty1212):

no that is not going to work

OpenStudy (misty1212):

we can complete the square maybe

OpenStudy (anonymous):

Can you help me complete the square?

OpenStudy (misty1212):

sure

OpenStudy (misty1212):

it is easier then the quadratic formula in this example\[k^2-24k+6=0\\ k^2-24k=-6\] is a start now what is half of \(24\)?

OpenStudy (anonymous):

(24/2)^2 = 12^2 = 144

OpenStudy (misty1212):

got so \[(k-12)^2=-6+144\\ (k-12)^2=138\]

OpenStudy (misty1212):

then \[k-12=\pm\sqrt{138}\\ k=12\pm\sqrt{138}\] and you are done

OpenStudy (anonymous):

Thank you!!

OpenStudy (misty1212):

\[\color\magenta\heartsuit\]